OpenStudy (anonymous):

16 ^-1/4 over 4 ^1/2

OpenStudy (anonymous):

\[(1/\sqrt[4]{16})/\sqrt{4}\]

OpenStudy (anonymous):

=(1/2)/2 =1/4

OpenStudy (anonymous):

the possible answers are 4,1,2,-1

OpenStudy (anonymous):

can you draw the question? or write the equation? please XD

OpenStudy (anonymous):